# LInK: Learning Joint Representations of Design and Performance Spaces through Contrastive Learning for Mechanism Synthesis

### v1

@@ 3.2.2 Batch BFGS Optimization on GPU
The BFGS approach for gradient-based optimization has been established as a robust method for unconstrained non-linear gradient-based optimization. However, there are two important matters that we must address before being able to apply this method to our problem. First, our problem is not unconstrained. This is because if the initial positions of the mechanism being optimized are moved to a position that leads to the term in the inverse cosine seen in equation 11, becoming greater than 1 or less than -1, we will have a locking/infeasible mechanism. As such, instead of using a simple line search for the optimization steps in BFGS, we use a heuristic approach to prevent infeasible and locking configurations. We thus use the common Wolfe conditions (Wolfe, 1971) for inexact line search to find the step size in each iteration of optimization; however, on top of the Wolfe conditions, we also check that a given stepsize does not lead to infeasible mechanisms by checking if the simulation has resulted in any NaN values.
The traditional approach of performing BFGS optimization individually for each mechanism is inefficient and time-consuming, particularly when leveraging CPU computation. To address this and take advantage of GPU acceleration, we have developed a new batch BFGS scheme that vectorizes the BFGS Hessian updates and line searches across multiple mechanisms simultaneously. This adjustment allows us to process all optimizations in parallel, significantly speeding up the refinement of candidates and speeding up the generation of optimal solutions. For the sake of brevity, further details on how this is done are omitted here; interested readers are referred to our publicly available code for the detailed implementation.
@@ 3.3 Implementation Details of the Optimization Framework
Our methodology begins by precomputing embeddings for 10 million mechanisms from the LINKS dataset using our contrastive learning model tailored for mechanisms. For any provided target curve, we normalize it as outlined earlier, including a smoothing step for hand-drawn or rough curves. This smoothing involves a fast Fourier transform to retain only the first seven frequencies, which we use to reconstruct and compute the curve’s embedding. All further optimization processes utilize the original, unsmoothed curves. This step is only done for computing the embedding of the target curve; the other optimization steps use the original input curves directly.
Given the embedding for the target curve, we then measure the cosine distance between the target embedding and all of the precomputed embeddings of the mechanisms in our dataset and pick the top 500 most similar (i.e., lowest cosine distance) mechanisms for the optimization. As mentioned earlier, we then perform 10 steps of batch BFGS on all of the 500 mechanisms and the target curves. To do this, we normalize the target curve and the paths generated by the candidate mechanisms, then find the optimal orientation of the target for the curves the candidate mechanisms produce. We then perform a brute-force grid search with 200 equally spaced rotation angles (equally spaced from 00 to $2\pi$), and identify the optimal rotation that minimizes the Chamfer distance. This rotation is identified for each mechanism in the candidate set. The target curves are then rotated to match the candidate paths as well as possible. Then we perform the BFGS with curves that have been oriented properly. Once the initial 10 steps of BFGS are performed, we then pick the top 50 best-performing at this stage of optimization and perform an additional 150 steps of BFGS on them. We call this ‘deep optimization’ framework for path synthesis the ‘Learning-accelerated Inverse Kinematics (LInK)‘. We look at LInK’s performance and efficiency in the following sections.
@@ 3.4 Manufacturability Considerations
So far we have only focused on the topic of kinematics and path synthesis. However, not all mechanisms that can be kinematically solved can additionally be manufactured. These mechanisms involve many linkages that move in space that have to be staked in layers, and these linkages can collide with each other; since each joint can be attached to multiple linkages, the joints may have to exist in many layers, and can therefore collide with linkages in each layer. This issue can be overcome with exotic designs for joints that enable circumventing collisions, but that would be costly and difficult to achieve in practice. As such, we must establish if any given mechanism is manufacturable with conventional parts. We can thus formulate an optimization problem to obtain the optimal configuration for a given linkage mechanism such that the total number of layers needed to manufacture a mechanism is minimized.
Each linkage will have a $z$ value, which we denote with $z_{i}$ for the $i$-th linkage in the set of linkages $L$. Each joint has an attachment to some number of linkages and collides with a set number of linkages as well. For each joint $j$ in the set of all joints $J$, we denote the set of linkages attached to it as $A_{j}$ and the set of linkages colliding with the joint as $C_{j}$ for the $j$-th joint. Finally, each linkage is in collision with some other linkages. We denote the set of linkages colliding with the $i$-th linkage as $O_{i}$, which excludes self-collision. Finally, a subset of $J$ denoted $G\subset J$ is the set of grounded joints. Our objective is to minimize the maximum of $z_{i}$ for all linkages. To obtain $O_{i}$, $A_{j}$ and $C_{j}$, we simulate the mechanism and find the colliding geometries. Given the above notation, we have the following design variables (the set $X$ contains all design variables):
@@
$M\text{ which will represent the maximum value amongst }z_{i}$
@@
This optimization simulation will yield z values which will be integers ranging from 0 to a maximum of $|L|-1$, which would be equivalent to having one layer per linkage. It is also important to note that the collisions for fixed joints are not considered. This is because each linkage can be grounded in space without the need for a joint pass through the mechanism. This formulation lends itself very well to mixed-integer linear programming (MILP) algorithms. In this case, we use Gurobi to solve the problem and determine if any given mechanism yields a feasible MILP problem; if so, what the optimal configuration for that given mechanism is. Note that since all of the operations in our approach happen in batches, at every step, we obtain many candidates, and as such we easily check the manufacturability of each mechanism from best to worst performing until the most manufacturable mechanism is found. This is an important aspect of the process that prior methods have ignored; few check for manufacturability at all. In our work from here on to the end, all mechanisms we identify as solutions are manufacturable; the above MILP is feasible for the mechanisms that are shown and discussed in this paper. For more details on this aspect of the work and visualizations of linkage assemblies please see Appendix C.
4 Results & Discussion
@@
In this section, we conduct a few different sets of experiments using our model and compare the results of our approach to the state of the art. Before discussing the details of the experiment, it is important to first discuss our experimental settings and the evaluation metrics we use to evaluate both our method and other approaches.
4.1 Evaluation Metrics & Experiment Details
@@
As we discussed before in path synthesis the primary aim is to come up with mechanisms that allow for the tracing of a desired target path. As such to measure the performance of any model for this task we need to look at metrics that allow for comparing two given paths purely based on their shape/path. Here we track two highly correlated metrics but each provides slightly different insight. These metrics are the Chamfer distance and ordered distance as discussed in Equation 12 and Equation 13 respectively. We also discussed the trade-offs of each metric with regards to comparing paths in Section 3.2.1. However, as far as comparison
@@
Finally, it is important to note that in all of our tests, we resample all curves, both target and traced curves by the mechanisms, by interpolating the curves to 2000 equidistant points. This is necessary to remove timing bias in the curves traced by mechanisms since the motion of the mechanism is not at a constant velocity. Removing timing is needed since the objective of path synthesis is to trace a path without any prescribed kinematics such as speed and acceleration. Now that the metrics and how we measure them are established we will conduct some tests to compare our method to prior works and finally, we will introduce two new benchmarks for this problem to establish a concrete benchmark for future works to try to handle.
@@ 4.2 Path Synthesis Case Studies
Figure 4: This figure shows the 8 test curves used in prior works (Left) and 24 random samples from the LINKS test subset (Middle) and the alphabet test curves (Right).
We will first test our model on three sets of problems and report the results of our approach on each of these test cases. Currently, the existing literature lacks standardized test sets. As such, it is hard for us to compare our model against the state of the art as these methods have limited code availability, and replicating them can be complicated. Despite this, a specific set of test curves have been proposed by Pan et al. (2023), which have been evaluated using multiple methods. However, this set is very small (i.e., only 8 curves), and is not well suited for rigorous testing of path synthesis methods such as ours. However, we use this dataset as it provides an avenue for comparison to the state of the art.
Table 1: This table presents the quantitative results of our model on the three different test sets we use to evaluate our model. For each test case, we run our model 5 times and report the best solution found within the 5 runs. The numbers following $\pm$ are the standard deviation across the samples in each test case. Note that although we normalize the LINKS test curves and the alphabet test curves we use the original scaling on the 8 test curves as they were measured in prior works at that scale.
@@
LINKS Test Subset Curves Generated By Random Planar Mechanisms 0.0057 $\pm$ 0.0039 0.0016 $\pm$ 0.0042
Popular 8 Curves Test Hand-Drawn Curves 0.034 $\pm$ 0.029 0.015 $\pm$ 0.021
Alphabet Benchmark Curves Mimicking Capital Letters of The English Alphabet 0.051 $\pm$ 0.041 0.229 $\pm$ 0.671
@@
Moving beyond, we also introduce two sets of benchmark problems, each of which provides a different insight into the performance of any method on this problem. First, we use in our approach 200 curves traced by a subset of the LINKS dataset Heyrani Nobari et al. (2022), which we set aside during training for the purpose of testing our method. These curves represent problems that we call ‘in-distribution,‘ as these curves are part of the same distribution of data used for training the model. However, it is important to test any method upon curves that are very different from the distribution of the data and provide an additional challenge by being particularly exotic shapes; curves which planar linkage mechanisms would be ordinarily considered capable of producing. In line with this, we introduce a rather challenging problem: tracing the capital letters in the English alphabet. To do this, we provide 26 continuous curves that align with the shapes of the capital letters in the English alphabet. This set includes shapes that are drastically different from the typical curves in the dataset, hence giving us an ‘out-of-distribution’ test case. Furthermore, these curves are riddled with sharp corners and features that are particularly difficult for planar linkage mechanisms to handle.
@@
Figure 5: This figure shows the best results from LInK on the 8 test curves used in prior works and the overlayed resulting path traced by the mechanism and the target. As it is visible here LInK has matched the target curves effectively without much challenge. For visualization of the underlying mechanisms please see Appendix C.
@@
Figure 6: This figure shows 20 random results from LInK on the larger LINKS test subset. We do not visualize mechanisms and only visualize the curve traced by the mechanisms to save space. As it can be seen in almost all cases the solution found is nearly identical to the target curve, which shows that when it comes to in-distribution target curves LInK performs exceptionally well.
@@
Figure 7: Right: This figure shows results from LInK on the larger alphabet test set. As can be seen, the paths produced by the generated mechanisms do not capture the alphabet shapes as well as we see in the other test sets. This shows that despite LInK performing well even on this test this test set presents a more challenging benchmark for path synthesis and can serve as a good benchmark for future methods. The ordered distance value for each curve is displayed above the curves for each alphabet path. Left: Showing the mechanism that traces B and the feasible region highlighted in blue where the highlighted joint can move without causing a singularity. This shows the non-linear and discontinuous nature of the problem and highlights the challenge of performing kinematic synthesis. Note that this is only assuming all other joints keep the same positions. Changing the position of any joint could change the feasible region for all other joints making the problem very challenging from the perspective of conventional optimization.
@@
Each of the three cases discussed above is visualized in figure 4, which displays all of the curves in the alphabet and the test curves proposed by Pan et al. (2023), while showing a random subset of the dataset test curves. We run our model for all three test cases and display the results visually in figures 7 and 5 for the alphabet and 8 test curves respectively. Finally, we show a few random example solutions from the LINKS test set in figure 6. As it can be seen when it comes to the 8 test curves our model performs exceptionally well and as we will see later when compared to the state of the art our approach significantly outperforms prior methods. Furthermore, we see that the model’s performance on the in-distribution test case is also very good, which should be expected as these samples are similar to the training curves. However, we can see that the curves produced for the alphabet test case do not follow a similar trend. Although for some letters the model has performed well, for many of the complex curves the results are under-whelming and do not trace the curves as expected. This test set, therefore, serves as a great benchmark for future methods as it presents a more demanding challenge, given the other challenges we test our model seem to be handled relatively easily.
@@
Beyond the qualitative results discussed so far, we measure the performance of our model on each of the test sets and report both the ordered distance and Chamfer distance in table 1. As it can be seen as we saw earlier the qualitative observations we had are indeed confirmed with both metrics, with the model performing well on the in-distribution test and the 8 test curves, while performing notably worse on the challenging alphabet curves (see the ordered distance of the alphabet test).
@@ 4.3 Comparison To State of Art Methods
Table 2: This table presents the performance metrics for our method LInK compared to the state-of-the-art deep learning model GCP-HOLO proposed by Fogelson et al. (2023) and optimization-based methods MINLP and MICP proposed by Pan et al. (2023). The values reported here are ordered distance in Eqn.13. The values for GCP-HOLO, MICP, and MINLP are reported based on the measurements in the original papers Fogelson et al. (2023); Pan et al. (2023). We observe that on average, LInK has 61X less error than GCP-HOLO, while being 33X faster (shown in Table 3)
@@
[Uncaptioned image] 0.0094 2.53 0.079 0.026 2.73 0.71 2.58
[Uncaptioned image] 0.0086 0.42 0.012 0.012 0.32 1.25 1.37
[Uncaptioned image] 0.013 0.21 0.015 0.013 0.44 0.70 0.77
[Uncaptioned image] 0.0041 2.01 0.23 0.061 1.58 1.39 2.52
[Uncaptioned image] 0.018 0.85 0.28 0.024 0.32 1.25 1.37
[Uncaptioned image] 0.0039 0.34 0.0063 0.0056 1.07 0.77 1.28
[Uncaptioned image] 0.0043 0.16 0.010 0.010 0.87 1.36 1.36
[Uncaptioned image] 0.060 0.45 0.090 0.072 0.38 0.36 0.32
Mean 0.015 0.87 0.089 0.029 0.92 0.97 1.45
In their work, Pan et al. (2023) proposed 8 curves for which they used MICP to perform path synthesis and reported their results. Since then, the current state-of-the-art method in path synthesis using deep learning (i.e., GCP-HOLO Fogelson et al. (2023)), has evaluated their method on this set as well. As such, we will test our method on this set of target paths as well, and compare our performance to the state of the art based on these paths.
In Table 2, we present the results of our approach for each of the 8 curves: both GCP-HOLO and MICP are used, and as we can see, LInK significantly out-performs these approaches by an order of magnitude demonstrating more than 94% improvement over the state of the art. One of the main explanations for the shortcomings of the approaches mentioned in this study is that these approaches are limited to smaller mechanisms, due to the high computational cost of dealing with increasingly larger mechanisms with these models. Despite limiting their size, these models are still significantly slower than our proposed framework. We discuss the efficiency and inference speed in later sections. Regardless, we run LInK on limited-sized datasets to demonstrate that LInK outperforms the state-of-the-art even when limited to smaller mechanisms. For these runs, we only conduct the contrastive search on the subset of mechanisms that are smaller than a certain threshold (See LInK(8) through LInK(16) in table 2).
4.4 Inference Speed & Scalability

### v2

@@ 3.2.2 Batch BFGS Optimization on GPU
The BFGS approach for gradient-based optimization has been established as a robust method for unconstrained nonlinear gradient-based optimization. However, there are two important matters that we must address before being able to apply this method to our problem. First, our problem is not unconstrained. This is because if the initial positions of the mechanism being optimized are moved to a position that leads to the term in the inverse cosine seen in equation 11, becoming greater than 1 or less than -1, we will have a locking/infeasible mechanism. As such, instead of using a simple line search for the optimization steps in BFGS, we use a heuristic approach to prevent infeasible and locking configurations. We thus use the common Wolfe conditions (Wolfe, 1971) for inexact line search to find the step size in each iteration of optimization; however, on top of the Wolfe conditions, we also check that a given step size does not lead to infeasible mechanisms by checking if the simulation has resulted in any NaN values.
The traditional approach of performing BFGS optimization individually for each mechanism is inefficient and time-consuming, particularly when leveraging CPU computation. To address this and take advantage of GPU acceleration, we have developed a new batch BFGS scheme that vectorizes the BFGS Hessian updates and line searches across multiple mechanisms simultaneously. This adjustment allows us to process all optimizations in parallel, significantly speeding up the refinement of candidates and the generation of optimal solutions. For the sake of brevity, further details on how this is done are omitted here; interested readers are referred to our publicly available code for the detailed implementation.
@@ 3.3 Implementation Details of the Optimization Framework
Our methodology begins by precomputing embeddings for 10 million mechanisms from the LINKS dataset using our contrastive learning model tailored for mechanisms. For any provided target curve, we normalize it as outlined earlier, including a smoothing step for hand-drawn or rough curves. This smoothing involves a fast Fourier transform to retain only the first seven frequencies, which we use to reconstruct the curve and compute the curve’s embedding. All further optimization processes use the original, unsmoothed curves. This step is only done for computing the embedding of the target curve; the other optimization steps use the original input curves directly.
Given the embedding for the target curve, we then measure the cosine distance between the target embedding and all of the precomputed embeddings of the mechanisms in our dataset and pick the top 500 most similar (i.e., lowest cosine distance) mechanisms for the optimization. As mentioned earlier, we then perform 10 steps of batch BFGS on all of the 500 mechanisms and the target curves. To do this, we normalize the target curve and the paths generated by the candidate mechanisms, then find the optimal orientation of the target for the curves the candidate mechanisms produce. We then perform a brute-force grid search with 200 equally-spaced rotation angles (from 00 to $2\pi$), and identify the optimal rotation that minimizes the Chamfer distance. This rotation is identified for each mechanism in the candidate set. The target curve is rotated to match the candidate paths as well as possible. Then we perform the BFGS with curves that have been oriented properly. Once the initial 10 steps of BFGS are performed, we then pick the top 50 best-performing at this stage of optimization and perform an additional 150 steps of BFGS on them. We call this deep optimization framework for path synthesis the Learning-accelerated Inverse Kinematics (LInK). We look at LInK’s performance and efficiency in the following sections.
Please note that the number of candidates and optimization steps at each stage are choices made empirically. We will demonstrate that these choices yield highly accurate solutions in rapid time compared to prior works. This approach strikes a balance between speed and accuracy. The optimization process can be made more computationally intensive if desired, potentially allowing for further improvements in the final solution. However, our current parameters have proven effective in achieving excellent results within reasonable time constraints.
@@ 3.4 Manufacturability Considerations
So far, we have only focused on the topics of kinematics and path synthesis, however, not all mechanisms that can be kinematically solved can additionally be manufactured. In a mechanism, linkages move freely through space, and therefore are typically stacked in layers to avoid collision. Linkages, or multi-layer joints, that occupy the same space as other linkages must be avoided, which is often resolved by adding more layers to a mechanism, increasing its complexity and decreasing its feasibility. This issue can also be overcome with exotic designs for joints that enable circumventing collisions, but that would be costly and difficult to achieve in practice. As such, we must establish if any given mechanism is manufacturable with conventional parts. We can thus formulate an optimization problem to obtain the optimal configuration for a given linkage mechanism such that the total number of layers needed to manufacture a mechanism is minimized.
Each linkage will have a $z$ value, which we denote with $z_{i}$ for the $i$-th linkage in the set of linkages $L$. Each joint has an attachment to some number of linkages and collides with a set number of linkages as well. For each joint $j$ in the set of all joints $J$, we denote the set of linkages attached to it as $A_{j}$, and the set of linkages colliding with the joint as $C_{j}$ for the $j$-th joint. Finally, each linkage is in collision with some other linkages. We denote the set of linkages colliding with the $i$-th linkage as $O_{i}$, which excludes self-collision. Finally, a subset of $J$ denoted $G\subset J$ is the set of grounded joints. Our objective is to minimize the maximum of $z_{i}$ for all linkages. To obtain $O_{i}$, $A_{j}$ and $C_{j}$, we simulate the mechanism and find the colliding geometries. Given the above notation, we have the following design variables (the set $X$ contains all design variables):
@@
$M\text{ which will represent the maximum value amongst }z_{i}.$
@@
This optimization simulation will yield z values which will be integers ranging from 0 to a maximum of $|L|-1$, which would be equivalent to having one layer per linkage. It is also important to note that the collisions for fixed joints are not considered. This is because each linkage can be grounded in space without the need for a joint to pass through the mechanism. This formulation lends itself very well to mixed-integer linear programming (MILP) algorithms. In this case, we use Gurobi to solve the problem and determine if any given mechanism yields a feasible MILP problem; and if so, determine what the optimal configuration for that given mechanism is. Note that since all of the operations in our approach happen in batches, at every step, we obtain many candidates, and as such we easily check the manufacturability of each mechanism from best to worst performing until the most manufacturable mechanism is found. This is an important aspect of the process that prior methods have ignored; few check for manufacturability at all. In our work from here on to the end, all mechanisms we identify as solutions are manufacturable; the above MILP is feasible for the mechanisms that are shown and discussed in this paper. For more details on this aspect of the work and visualizations of linkage assemblies please see Appendix D.
4 Results and Discussion
@@
In this section, we share the results of a few different sets of experiments using our model and compare the results of our approach to the state of the art. Before discussing the details of the experiment, it is important to first discuss our experimental settings and the evaluation metrics we use to evaluate both our method and other approaches.
4.1 Evaluation Metrics and Experiment Details
@@
As we discussed before in path synthesis, the primary aim is to come up with mechanisms that allow for the tracing of a desired target path. As such, to measure the performance of any model for this task we need to look at metrics that allow for comparing two given paths purely based on their shape/path. Here we track two highly correlated metrics but, each provides slightly different insight. These metrics are the Chamfer distance and ordered distance as discussed in Equation 12 and Equation 13 respectively. We discussed the trade-offs of each metric with regards to comparing paths in Section 3.2.1.
@@
It is also important to note that in our tests we resample all target and traced curves (by the mechanisms) by interpolating the curves to 2000 equidistant points. This is necessary to remove timing bias in the curves traced by mechanisms, since the motion of the mechanism is not at a constant velocity. Removing timing is needed since the objective of path synthesis is to trace a path without any prescribed kinematics such as speed and acceleration. Now that the metrics, and how we measure them are established, we will conduct some tests to compare our method to prior works. Finally, we will introduce two new benchmarks for this problem to establish a concrete benchmark for future works.
@@ 4.2 Path Synthesis Case Studies
Figure 6: This figure shows the 8 test curves used in prior works (Left), 24 random samples from the LINKS test subset (Middle), and the alphabet test curves (Right).
We will first test our model on three sets of problems and report the results of our approach on each of these test cases. Currently, the existing literature lacks standardized test sets. As such, it is hard for us to compare our model against the state of the art, as these methods have limited code availability and replicating them can be complicated. Despite this, a specific set of test curves have been proposed by Pan et al. (2023), which have been evaluated using multiple methods. However, this set is very small (i.e., only 8 curves), and is not well suited for rigorous testing of path synthesis methods such as ours. However, we use this dataset, as it provides an avenue for comparison to the state of the art.
Table 1: This table presents the quantitative results of our model on the three different test sets we used to evaluate our model. For each test case, we run our model 5 times and report the best solution found within the 5 runs. The numbers following $\pm$ are the standard deviation across the samples in each test case. Note that although we normalize the LINKS and alphabet test curves, we use the original scaling on the 8 test curves, as they were measured in prior works at that scale.
@@
LINKS Test Subset Curves Generated By Random Planar Mechanisms 0.0059 $\pm$ 0.0029 0.0018 $\pm$ 0.0043
Popular 8 Curves Test Hand-Drawn Curves 0.026 $\pm$ 0.011 0.0106 $\pm$ 0.0083
Alphabet Benchmark Curves Mimicking Capital Letters of The English Alphabet 0.0825 $\pm$ 0.2301 0.179 $\pm$ 0.737
@@
Moving beyond, we also introduce two sets of benchmark problems, each of which provides a different insight into the performance of any method on this problem. First, we use in our approach 200 curves traced by a subset of the LINKS dataset Heyrani Nobari et al. (2022), which we set aside during training for the purpose of testing our method. These curves represent problems that we call in-distribution, as these curves are part of the same distribution of data used for training the model. However, it is important to test any method upon curves that are very different from the distribution of the data, to provide an additional challenge by being particularly exotic shapes; curves which planar linkage mechanisms would be ordinarily considered capable of producing. In line with this, we introduce a rather challenging problem: tracing the capital letters in the English alphabet. To do this, we provide 26 continuous curves that align with the shapes of the capital letters in the English alphabet. This set includes shapes that are drastically different from the typical curves in the dataset, hence giving us an out-of-distribution test case. Furthermore, these curves are riddled with sharp corners and features that are particularly difficult for planar linkage mechanisms to handle.
@@
Figure 7: This figure shows the best results from LInK on the 8 test curves used in prior works. The overlayed resulting path traced by the mechanism and the target is shown in red, and the target curve is shown in black. As it is visible here, LInK has matched the target curves effectively without much challenge. For visualization of the underlying mechanisms please see Appendix D.
@@
Figure 8: This figure shows 20 random results from LInK on the larger LINKS test subset. We do not visualize mechanisms and only visualize the curve traced by the mechanisms to save space. As it can be seen in almost all cases, the solution found is nearly identical to the target curve, which shows that when it comes to in-distribution target curves, LInK performs exceptionally well.
@@
Figure 9: This figure shows results from LInK on the larger alphabet test set. As can be seen, the paths produced by the generated mechanisms do not capture the alphabet shapes as well as we see in the other test sets. This shows that despite LInK performing well even on this test, this test set presents a more challenging benchmark for path synthesis and can serve as a good benchmark for future methods. The ordered distance value for each curve is displayed above the curves for each alphabet path.
@@
Each of the three cases discussed above are visualized in Figure 6, which displays all of the curves in the alphabet and the test curves proposed by Pan et al. (2023), while showing a random subset of the dataset test curves. We run our model for all three test cases and display the results visually in figures 9 and 7 for the alphabet and 8 test curves respectively. Finally, we show a few random example solutions from the LINKS test set in Figure 8. As it can be seen when it comes to the 8 test curves our model performs exceptionally well and as we will see later when compared to the state of the art our approach significantly outperforms prior methods. Furthermore, we see that the model’s performance on the in-distribution test case is also very good, which should be expected as these samples are similar to the training curves. However, we can see that the curves produced for the alphabet test case do not follow a similar trend. Although for some letters the model has performed well, for many of the complex curves the results are under-whelming and do not trace the curves as expected. This test set, therefore, serves as a great benchmark for future methods as it presents a more demanding challenge.
@@
Beyond the qualitative results discussed so far, we measure the performance of our model on each of the test sets and report both the ordered distance and Chamfer distance in Table 1. As can be seen, the qualitative observations are confirmed with both metrics, with the model performing well on the in-distribution test and the 8 test curves, while performing notably worse on the challenging alphabet curves (see the ordered distance of the alphabet test).
@@ 4.3 Comparison To State of Art Methods
Table 2: This table presents the performance metrics for LInK compared to the state-of-the-art deep learning model, GCP-HOLO, proposed by Fogelson et al. (2023) and optimization-based methods, MINLP and MICP, proposed by Pan et al. (2023). The values reported here are ordered distance in Equation 13. The values for GCP-HOLO, MICP, and MINLP are reported based on the measurements in the original papers Fogelson et al. (2023); Pan et al. (2023). The numbers in brackets after LInK indicate the maximum number of joints LInK was allowed to use (The full LInK method can use up to 20 joints). We observe that on average, LInK has 61 times less error than GCP-HOLO, while being 150 times faster (shown in Table 3)
@@
[Uncaptioned image] 0.0144 0.1638 0.0337 0.0144 2.73 0.71 2.58
[Uncaptioned image] 0.0056 0.0257 0.0101 0.0059 0.32 1.25 1.37
[Uncaptioned image] 0.0106 0.024 0.0134 0.0106 0.44 0.70 0.77
[Uncaptioned image] 0.0039 0.0881 0.0121 0.0107 1.58 1.39 2.52
[Uncaptioned image] 0.0166 0.6052 0.022 0.0166 0.32 1.25 1.37
[Uncaptioned image] 0.0022 0.0155 0.0036 0.0022 1.07 0.77 1.28
[Uncaptioned image] 0.0034 0.0048 0.0048 0.0037 0.87 1.36 1.36
[Uncaptioned image] 0.0282 0.1265 0.0633 0.0282 0.38 0.36 0.32
Mean 0.0106 0.1317 0.0204 0.0115 0.92 0.97 1.45
In their work, Pan et al. (2023) proposed 8 curves for which they used MICP to perform path synthesis and reported their results. Since then, the current state-of-the-art method in path synthesis using deep learning (i.e., GCP-HOLO Fogelson et al. (2023)), has evaluated their method on this set. As such, we will test our method on this set of target paths as well, and compare our performance to the state of the art based on these paths.
In Table 2, we present the results of our approach for each of the 8 curves: both GCP-HOLO and MICP are used, and as we can see, LInK significantly outperforms these approaches by an order of magnitude demonstrating more than a 94% improvement over the state of the art. One of the main explanations for the shortcomings of the approaches mentioned in this study is that these approaches are limited to smaller mechanisms, due to the high computational cost of dealing with increasingly larger mechanisms with these models. Despite limiting their size, these models are still significantly slower than our proposed framework. We discuss the efficiency and inference speed in later sections. Regardless, we run LInK on limited-sized datasets to demonstrate that LInK outperforms the state of the art, even when limited to smaller mechanisms. To do this we run LInK with a limit of 7 to 20 joints for each curve and show these results in Figure 10. For these runs, we only conduct the contrastive search on the subset of mechanisms that have fewer than the indicated joints (See LInK(7) through LInK(15) in Table 2).
4.4 Inference Speed and Scalability